# A Centrality-independent Framework for Revealing Genuine Higher-Order Cumulants in Heavy-Ion Collisions

### v1

@@ 3.6 Uncertainty Estimation
In this step, we evaluate the uncertainties of the correction method through two distinct components: statistical uncertainty and systematic uncertainty. To estimate statistical uncertainty, we applied the Bootstrap resampling method to obtain results. For systematic uncertainty, we systematically apply several sources include varying the tightness of the constraint conditions, implementing overall upward and downward shifts for the $C_{1}$ values, and adjusting the upper and lower boundaries applied to the overall distribution. The final uncertainty estimation combines both components. So our final results with uncertainties are shown in Figure 11.
@@ 4 Summary
Looking ahead, this analysis framework has broad applicability and is particularly significant in low-energy collisions, where the impacts of initial volume fluctuations and centrality resolution are more pronounced, potentially obscure physical signals. It can be extended to study net-proton number fluctuations and other particle multiplicity fluctuations. The framework can be applied to the study of correlations such as baryon-strangeness ($B-S$) and baryon-charge ($B-Q$), as well as higher-order measurements. In the realistic data analysis with the presence of detector efficiency effects, this method enables the extraction of cumulants that are inherently unaffected by initial volume fluctuations, with subsequent efficiency corrections applied to these results. This approach allows us to extract genuine physical signals, thereby paving the way for probing the intrinsic thermodynamic properties of the hot dense QCD medium through event-by-event fluctuations. By effectively circumventing volume fluctuation effects, our method provides more reliable experimental measurements for exploring the QCD phase diagram [25, 26, 27].
@@ Acknowledgements
We extend our gratitude to Dr. Nu Xu for useful discussions. This work is supported in part by the National Key Research and Development Program of China under contract Nos. 2022YFA1604900 and the National Natural Science Foundation of China under Grant No. 12447102.

### v2

@@ 3.6 Uncertainty Estimation
We evaluate the uncertainties associated with the correction method by considering two distinct components: statistical uncertainty and systematic uncertainty. The statistical uncertainty is estimated using the Bootstrap resampling method. For systematic uncertainty, we incorporate multiple sources including variations in constraint condition tightness and adjustments to the upper and lower boundaries of the overall distribution. The final uncertainty estimation combines both components, and our comprehensive results with uncertainties are presented in Fig. 11.
3.7 Method Validation with Different Centrality Resolutions
Based on UrQMD model, we conducted a comprehensive evaluation of the method’s performance under various centrality resolutions in Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=3.5$ GeV. Specifically, we examined four distinct centrality determination methods: RefMult3 without scaling, RefMult3 with scaling, RefSpect3 without scaling, and RefSpect3 with scaling. Here, RefMult3 refers to the charged particle multiplicity within the $\eta$ range (0, 2.4), while RefSpect3 denotes charged particle multiplicity within the $\eta$ range (3.3, 6.0).
Figure 12: Charged particle multiplicity distributions for different centrality determination in Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=3.5$ GeV from UrQMD model.
Figure 13: Proton cumulants and cumulant ratios calculated with the method based on different centrality resolutions in Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=3.5$ GeV from UrQMD model. A 2.5% centrality interval is used in this test.
The scaling procedure accounts for detector efficiency effects by applying scaling factors of 0.7 and 0.8 to the charged particle multiplicities for RefMult3 and RefSpect3, respectively. This scaling approach effectively reduces the resolution by compensating for detector inefficiencies. The charged particle multiplicity distributions corresponding to these four resolution configurations are presented in Fig. 12. The resolution ranked in descending order: RefMult3 (without scaling), RefMult3 (with scaling), RefSpect3 (without scaling), RefSpect3 (with scaling). As demonstrated in Fig. 13, our approach produces consistent results across all four centrality resolution configurations. This provides compelling evidence for the robustness and reliability of the method.
@@ 4 Summary
Looking ahead, this analysis framework has broad applicability and is particularly significant in low-energy collisions, where the impacts of initial volume fluctuations and centrality resolution are more pronounced, potentially obscure physical signals. It can be extended to study net-proton number fluctuations and other particle multiplicity fluctuations. The framework can be applied to the study of correlations such as baryon-strangeness (BSB-S) and baryon-charge (BQB-Q), as well as higher-order measurements. In the realistic data analysis with the presence of detector efficiency effects, this method enables the extraction of cumulants that are inherently unaffected by initial volume fluctuations, with subsequent efficiency corrections applied to these results. This approach allows us to extract genuine physical signals, thereby paving the way for probing the intrinsic thermodynamic properties of the hot dense QCD medium through event-by-event fluctuations. By effectively circumventing volume fluctuation effects, our method provides more reliable experimental measurements for exploring the QCD phase diagram [Chen:2024aom, Bzdak:2019pkr, Luo:2022mtp].
@@ Acknowledgements
We extend our gratitude to Dr. Nu Xu for useful discussions. This work is supported in part by the National Key Research and Development Program of China under contract Nos. 2022YFA1604900 and the National Natural Science Foundation of China under Grant No. 12525509 and No. 12447102.